# Quantum Dynamic Programming

### v1

@@ IV.4 From Lemma 18 to Theorem 21
In Lemma 18, we have shown that the distance $\delta_{n}$ between the final state and the target state $\tau$ after $n$ QDP Grover search iterations follows the recursion relation Eq. (132). In this subsection, we demonstrate that for a reasonably small implementation error $\epsilon$, the sequence $\delta_{n}$ does not deviate from the sequence of distance obtained from ideal, errorless iterations. To do this, we set $\delta_{n+1}=h(\delta_{n})+\epsilon$, with the function
$h(x)=\operatorname{sech}\left((2L+1)\operatorname{arcsech}(x)\right)\ .$ (140)
Note that Eq. (140) gives distance recurrence relations for the exact nested fixed-point Grover search (Lemma 10), i.e. $\frac{1}{2}\lVert\tau-\psi_{n}\rVert_{1}=h(\frac{1}{2}\lVert\tau-\psi_{n-1}\rVert_{1})$ for $|{\psi_{n}}\rangle=\hat{\mathbf{\Gamma}}^{(\psi_{n-1})}_{L}|{\psi_{n-1}}\rangle$. By finding the range for fast spectral convergence (Def. 8), i.e. the range where its derivative $h^{\prime}(x)<r$ for some $r<1$ and $h(x)+\varepsilon<x$, we demonstrate that the resulting distance from $N$ iterations of the QDP implementation $\delta_{N}$ can still be arbitrarily close to that of the exact implementation.
@@
The function $h(x)$ defined as in Eq. (140) is monotonically increasing and convex in the range $x\in(0,1)$.
@@
$h^{\prime}(x)$ :-$\coloneq\frac{dh(x)}{dx}=\frac{\tilde{L}\operatorname{sech}\left(\tilde{L}\operatorname{arcsech}(x)\right)\tanh\left(\tilde{L}\operatorname{arcsech}(x)\right)}{x\sqrt{1-x^{2}}}>0\ ,:$ (141)
@@
$g(y)\coloneq\frac{\tilde{L}\operatorname{sech}\left(\tilde{L}y\right)\tanh\left(\tilde{L}y\right)}{\operatorname{sech}(y)\tanh(y)}\ ,$ (142)
@@
$h^{\prime\prime}(x)\coloneq\frac{d^{2}h(x)}{dx^{2}}=\frac{dg(y)}{dy}\frac{dy}{dx}=-\frac{g^{\prime}(y)}{x\sqrt{1-x^{2}}}\ ,$ (143)
@@
$g^{\prime}(y)=\frac{\tilde{L}}{y}\frac{\cosh^{2}(y)\sinh(\tilde{L}y)}{\sinh(y)\cosh^{2}(\tilde{L}y)}\left[\tilde{L}y\frac{2\operatorname{sech}^{2}(\tilde{L}y)-1}{\tanh(\tilde{L}y)}-y\frac{2\operatorname{sech}^{2}(y)-1}{\tanh(y)}\right]\ .$ (144)
@@
$f(y)\coloneq y\frac{2\operatorname{sech}^{2}(y)-1}{\tanh(y)}\ .$ (145)
@@
$x_{2}=1-\frac{1}{4(L^{2}+L)}\epsilon+\mathcal{O}(\epsilon^{2})\ .$ (146)
@@
$\tilde{L}\sqrt{2y}=\sqrt{2(y+\epsilon)}+\mathcal{O}(y^{3/2})$ (147)
@@
Consider the nested fixed-point Grover search with $N$ recursions and $2L$ partial reflections at each recursion step. Furthermore, assume that the distance between the initial state $\psi_{0}$ and the target state $\tau$ satisfies $\frac{1}{2}\lVert\tau-\psi_{0}\rVert_{1}\eqcolon\delta_{0}<\delta^{*}$, where $\delta^{*}$ is determined by $h^{\prime}(\delta^{*})=1$ with $h^{\prime}$ defined in Eq. (141). Then, QDP implementation of the algorithm can output a state $\rho_{N}$, whose distance to the target state $\tau$ is arbitrarily close to that of the $|{\psi_{N}}\rangle$, an output from the exact algorithm, i.e.
$\frac{1}{2}\lVert\tau-\rho_{N}\rVert_{1}\leq\frac{1}{2}\lVert\tau-\psi_{N}\rVert_{1}+\eta\ ,$ (148)
for any $\eta>0$. In addition, the success probability of the QDP implementation exceeds $1-p_{\mathrm{th}}$ for any $p_{\mathrm{th}}<\frac{N}{\log(L)}$. This implementation requires $L^{\mathcal{O}(N)}\eta^{-N}\log(p_{\rm th}^{-1})$ copies of $|{\psi_{0}}\rangle$ and a circuit of depth $\mathcal{O}(L^{2}N\eta^{-1})$.
@@
$\frac{1}{2}\lVert\tau-\psi_{N}\rVert_{1}=\tilde{\delta}_{N},\quad\frac{1}{2}\lVert\tau-\psi_{0}\rVert_{1}=\delta_{0}\ ,$ (149)
@@
$\frac{1}{2}\lVert\tau-\rho_{N}\rVert_{1}\leq\delta_{N}+\frac{\epsilon}{2L\pi}\ ,$ (150)
@@
$\delta_{N}<h^{N}(\delta_{0})+\sum_{n=0}^{N-1}r^{n}\epsilon=\tilde{\delta}_{N}+\frac{1-r^{N}}{1-r}\epsilon\ .$ (151)
@@
In the unfolding implementation, the circuit depth scales as $(2L+1)^{N}\sim\log(\delta_{N}^{-1})(1-\delta_{0}^{2})^{-1/2}$ with the Grover scaling factor $(1-\delta_{0}^{2})^{-1/2}$. However, for sampling-based Grover search algorithms, i.e. if copies of the initial state $\psi_{0}$ are used instead of the oracle $\hat{E}^{(\psi_{0})}_{\alpha}$, Ref. [14] proved that at least $\mathcal{O}(L^{2N})=\mathcal{O}((1-\delta_{0}^{2})^{-1})$ copies of the initial state are needed. The QDP implementation also obeys this scaling and the number of copies in Thm. 21 must scale faster than $\mathcal{O}(L^{2N})$.
V Oblivious Schmidt decomposition using double-bracket iterations
Oblivious Schmidt decomposition (OSD) is done by diagonalizing a reduced state $\rho_{0}^{(A)}=\operatorname{Tr}_{B}[\psi_{0}]$ of the given entangled state $\psi_{0}$, without learning the state itself. For diagonalization, we use a recently-proposed quantum algorithm [12], which we call double-bracket iterations, that adapts the continuous gradient flow for matrices for discrete gate-based quantum computing models.
V.1 Double-bracket recursions
In this section, we expound on the double-bracket iteration algorithm, which can be used for scenarios beyond OSD. A double-bracket iteration step with an instruction operator $\hat{H}_{n}$ is defined as
$\hat{H}_{n+1}=e^{s_{n}[\hat{D},\hat{H}_{n}]}\hat{H}_{n}e^{-s_{n}[\hat{D},\hat{H}_{n}]}\ .$ (152)
In other words, this is an example of the single memory-call recursion unitary, where the memory-call is of the linear map $\mathcal{N}(\hat{H})=-is_{n}[\hat{D},\hat{H}]$. Flow duration $s_{n}$ and diagonal operator $\hat{D}$ are not part of the quantum instructions and can be chosen obliviously to $\hat{H}_{n}$, or strategically with some a priori information on $\hat{H}_{n}$.
@@ V.1 Double-bracket recursions
$f\left(\hat{H}\right)\coloneq\lVert\hat{H}-\hat{D}\rVert_{2}^{2}\ .$ (153)
The generator $\frac{d}{ds_{n}}\hat{H}_{n+1}=[[\hat{D},\hat{H}_{n}],\hat{H}_{n}]$ is shown to be $-\mathrm{grad}f(\hat{H}_{n})$ for some Riemannian metric defined on the manifold of matrices isospectral to $\hat{H}_{0}$ [24]. Hence, the iteration effectively applies the gradient flow of $f$ and we expect $f$ to reduce over the iterations towards the minimum. This is in fact true for suitable choices of $s_{n}$ and $\hat{D}$.
@@
When $\hat{D}$ is a non-degenerate diagonal matrix such that $\hat{D}=\sum_{i}\mu_{i}|{i}\rangle\!\langle{i}|$ with $\mu_{i}>\mu_{j}$ for all $i>j$ and $s_{n}=\frac{1}{4\lVert\hat{H}_{0}\rVert_{2}\lVert\hat{D}\rVert_{2}}$, the recursion Eq. (152) has a unique stable fixed-point $\hat{H}_{\infty}=\sum_{i}\lambda_{i}|{i}\rangle\!\langle{i}|$, where $\lambda_{i}$ are eigenvalues of $\hat{H}_{0}$ arranged in a non-increasing order. Furthermore, this fixed-point is locally exponentially stable.
In particular, the off-diagonal element of $\hat{H}_{k}$ is suppressed after each step
$(\hat{H}_{k+1})_{ij}=\left[1-\frac{(\lambda_{i}-\lambda_{j})(\mu_{i}-\mu_{j})}{4\lVert\hat{H}_{0}\rVert_{2}\lVert\hat{D}\rVert_{2}}\right](\hat{H}_{k})_{ij}\ ,$ (154)
up to the first order of $(\hat{H}_{k})_{ij}$, where the suppression factor inside the square bracket is in the range $(0,1)$. Hence, it is possible to establish the exponential suppression of the distance.
@@
Let $\hat{H}_{\infty}$ be a fixed-point of the double-bracket iterations starting from a matrix $\hat{H}_{0}$ and let $M(\hat{H}_{0})$ be a manifold of matrices isospectral to $\hat{H}_{0}$. Then there exists $r<1$ and a neighbourhood $N$ of $\hat{H}_{\infty}$ in the manifold $M(\hat{H}_{0})$, such that
$\lVert\hat{H}_{k+1}-\hat{H}_{\infty}\rVert_{2}<r\lVert\hat{H}_{k}-\hat{H}_{\infty}\rVert_{2}\ ,$ (155)
for any $\hat{H}_{k}\in N$ and $\hat{H}_{k+1}$ obtained from an iteration from $\hat{H}_{k}$.
@@ V.2 Unfolding implementation for double-bracket iterations with black box queries
We now describe the unfolding implementation of double-bracket iterations, given access to the black box queries to the evolution $e^{\pm it\hat{H}_{0}}$ by the root operator $\hat{H}_{0}$, following Ref. [12]. To implement the memory-call $e^{i\mathcal{N}(\hat{H})}=e^{s_{n}[\hat{D}_{n},\hat{H}]}$ with queries $e^{\pm it\hat{H}_{0}}$, the group commutator approximation is used.
@@
$\left\lVert e^{-i\sqrt{s}\hat{B}}e^{-i\sqrt{s}\hat{A}}e^{i\sqrt{s}\hat{B}}e^{i\sqrt{s}\hat{A}}-e^{s[\hat{A},\hat{B}]}\right\rVert\leq s^{3/2}\left(\left\lVert[\hat{A},[\hat{A},\hat{B}]\right\rVert+\left\lVert[\hat{B},[\hat{B},\hat{A}]\right\rVert\right)\ .$ (156)
Setting $\hat{A}=\hat{H}$ and $\hat{B}=-\hat{D}$ we obtain an locally accurate approximation to the recursion step
$\hat{W}_{n}=\left(e^{i\sqrt{\frac{s_{n}}{M}}\hat{D}_{n}}e^{-i\sqrt{\frac{s_{n}}{M}}\hat{H}_{n}}e^{-i\sqrt{\frac{s_{n}}{M}}\hat{D}_{n}}e^{i\sqrt{\frac{s_{n}}{M}}\hat{H}_{n}}\right)^{M}$ (157)
that approximates $e^{s_{n}[\hat{D},\hat{H}_{n}]}$ with error $\mathcal{O}(s_{n}^{3/2}M^{-1/2})$ by making $2M$ queries to the evolution. Typically, we apply this recursion to a state
$\rho_{n}=\frac{\hat{H}_{n}+\lambda\mathmybb{1}}{\operatorname{Tr}\left[\hat{H}_{n}+\lambda\mathmybb{1}\right]}\ ,$ (158)
encoding the matrix $\hat{H}_{n}$. The convergence of the recursion $\rho_{k+1}=\hat{W}_{k}\rho_{k}\hat{W}_{k}^{\dagger}$ to the diagonal state $\rho_{\infty}$ has been established, despite the error arising from the group commutator approximation [12].
The evolution with respect to $\hat{H}_{n}$ needed for Eq. (157) is not directly accessible to us. However, the evolution is covariant, i.e. $e^{-it\hat{H}_{n}}=\hat{W}_{n-1}e^{-it\hat{H}_{n-1}}\hat{W}_{n-1}^{\dagger}$. Unfolding then follows from replacing all evolutions in Eq. (157) by the evolution query to the root state $e^{\pm it\hat{H}_{0}}$, dressed with previous recursion steps. Suppose that the gate count for $n$ iterations is $c_{n}$; then the next iteration uses $2Mc_{n}$ gates to make queries for the evolutions $e^{\pm it\hat{H}_{n+1}}$, implying $c_{n+1}\geq 2Mc_{n}$. The total circuit depth then scales exponentially as $\mathcal{O}((2M)^{N})$ with the number of iterations $N$.

### v2

@@ IV.4 From Lemma 19 to Theorem 22
In Lemma 19, we have shown that the distance $\delta_{n}$ between the final state and the target state $\tau$ after $n$ QDP Grover search iterations follows the recursion relation Eq. (S121). In this subsection, we demonstrate that for a reasonably small implementation error $\epsilon$, the sequence $\delta_{n}$ does not deviate from the sequence of distance obtained from ideal, errorless iterations. To do this, we set $\delta_{n+1}=h(\delta_{n})+\epsilon$, with the function
$h(x)=\operatorname{sech}\left((2L+1)\operatorname{arcsech}(x)\right)\ .$ (S128)
Note that Eq. (S128) gives distance recurrence relations for the exact nested fixed-point Grover search (Lemma 13), i.e. $\frac{1}{2}\lVert\tau-\psi_{n}\rVert_{1}=h(\frac{1}{2}\lVert\tau-\psi_{n-1}\rVert_{1})$ for $|{\psi_{n}}\rangle=\hat{\mathbf{\Gamma}}^{[\psi_{n-1}]}_{L}|{\psi_{n-1}}\rangle$. By finding the range for fast spectral convergence (Def. 8), i.e. the range where its derivative $h^{\prime}(x)<r$ for some $r<1$ and $h(x)+\varepsilon<x$, we demonstrate that the resulting distance from $N$ iterations of the QDP implementation $\delta_{N}$ can still be arbitrarily close to that of the exact implementation.
@@
The function $h(x)$ defined as in Eq. (S128) is monotonically increasing and convex in the range $x\in(0,1)$.
@@
$h^{\prime}(x)$ :-$\coloneq\frac{dh(x)}{dx}=\frac{\tilde{L}\operatorname{sech}\left(\tilde{L}\operatorname{arcsech}(x)\right)\tanh\left(\tilde{L}\operatorname{arcsech}(x)\right)}{x\sqrt{1-x^{2}}}>0\ ,:$ (S129)
@@
$g(y)\coloneq\frac{\tilde{L}\operatorname{sech}\left(\tilde{L}y\right)\tanh\left(\tilde{L}y\right)}{\operatorname{sech}(y)\tanh(y)}\ ,$ (S130)
@@
$h^{\prime\prime}(x)\coloneq\frac{d^{2}h(x)}{dx^{2}}=\frac{dg(y)}{dy}\frac{dy}{dx}=-\frac{g^{\prime}(y)}{x\sqrt{1-x^{2}}}\ ,$ (S131)
@@
$g^{\prime}(y)=\frac{\tilde{L}}{y}\frac{\cosh^{2}(y)\sinh(\tilde{L}y)}{\sinh(y)\cosh^{2}(\tilde{L}y)}\left[\tilde{L}y\frac{2\operatorname{sech}^{2}(\tilde{L}y)-1}{\tanh(\tilde{L}y)}-y\frac{2\operatorname{sech}^{2}(y)-1}{\tanh(y)}\right]\ .$ (S132)
@@
$f(y)\coloneq y\frac{2\operatorname{sech}^{2}(y)-1}{\tanh(y)}\ .$ (S133)
@@
$x_{2}=1-\frac{1}{4(L^{2}+L)}\epsilon+\mathcal{O}(\epsilon^{2})\ .$ (S134)
@@
$\tilde{L}\sqrt{2y}=\sqrt{2(y+\epsilon)}+\mathcal{O}(y^{3/2})$ (S135)
@@
Consider the nested fixed-point Grover search with $N$ recursions and $2L$ partial reflections at each recursion step. Furthermore, assume that the distance between the initial state $\psi_{0}$ and the target state $\tau$ satisfies $\frac{1}{2}\lVert\tau-\psi_{0}\rVert_{1}\eqcolon\delta_{0}<\delta^{*}$, where $\delta^{*}$ is determined by $h^{\prime}(\delta^{*})=1$ with $h^{\prime}$ defined in Eq. (S129). Then, QDP implementation of the algorithm can output a state $\rho_{N}$, whose distance to the target state $\tau$ is arbitrarily close to that of the $|{\psi_{N}}\rangle$, an output from the exact algorithm, i.e.
$\frac{1}{2}\lVert\tau-\rho_{N}\rVert_{1}\leq\frac{1}{2}\lVert\tau-\psi_{N}\rVert_{1}+\eta\ ,$ (S136)
for any $\eta>0$. In addition, the success probability of the QDP implementation exceeds $1-p_{\mathrm{th}}$ for any $p_{\mathrm{th}}\in(0,1)$. This implementation requires $L^{(2+\Delta)N}\eta^{-N}\log(p_{\rm th}^{-1})$ copies of $|{\psi_{0}}\rangle$ and a circuit of depth $\mathcal{O}(L^{2}N\eta^{-1})$ with some constant $\Delta>0$.
@@
$\frac{1}{2}\lVert\tau-\psi_{N}\rVert_{1}=\tilde{\delta}_{N},\quad\frac{1}{2}\lVert\tau-\psi_{0}\rVert_{1}=\delta_{0}\ ,$ (S137)
@@
$\frac{1}{2}\lVert\tau-\rho_{N}\rVert_{1}\leq\delta_{N}+\frac{\epsilon}{2L\pi}\ ,$ (S138)
@@
$\delta_{N}<h^{N}(\delta_{0})+\sum_{n=0}^{N-1}r^{n}\epsilon=\tilde{\delta}_{N}+\frac{1-r^{N}}{1-r}\epsilon\ .$ (S139)
@@
In the unfolding implementation, the circuit depth scales as $(2L+1)^{N}\sim\log(\delta_{N}^{-1})(1-\delta_{0}^{2})^{-1/2}$ with the Grover scaling factor $(1-\delta_{0}^{2})^{-1/2}$. However, for sampling-based Grover search algorithms, i.e. if copies of the initial state $\psi_{0}$ are used instead of the oracle $\hat{E}^{[\psi_{0}]}_{\alpha}$, Ref. [16] proved that at least $\mathcal{O}(L^{2N})=\mathcal{O}((1-\delta_{0}^{2})^{-1})$ copies of the initial state are needed. The QDP implementation also obeys this scaling, as can be seen in the exponent $(2+\Delta)N$ for the width scaling in $L$.
V Double bracket iterations: quantum imaginary-time evolution and oblivious Schmidt decomposition
Two example algorithms—quantum imaginary-time evolution and oblivious schmidt decomposition—can be considered as an ordered diagonalization task: from a generic pure state to the ground state in the Hamiltonian basis in the former case and from a generic reduced state to the diagonal state in computational basis in the latter case. In both cases, we use a recently-proposed quantum algorithm [6], which we call double-bracket iterations, that adapts the continuous gradient flow for matrices for discrete gate-based quantum computing models.
V.1 Double-bracket iterations
In this section, we expound on the double-bracket iteration algorithm, which can be used for scenarios beyond quantum imaginary-time evolution or oblivious Schmidt decomposition. A double-bracket iteration step with an instruction operator $\hat{P}_{n}$ is defined as
$\hat{P}_{n+1}=e^{s_{n}[\hat{D},\hat{P}_{n}]}\hat{P}_{n}e^{-s_{n}[\hat{D},\hat{P}_{n}]}\ .$ (S140)
In other words, this is a single memory-call recursion unitary, where the memory-call is defined by the linear Hermitian-preserving map $\mathcal{N}(\hat{P})=-is_{n}[\hat{D},\hat{P}]$. Flow duration $s_{n}$ and diagonal operator $\hat{D}$ are not part of the quantum instructions and can be chosen obliviously to $\hat{P}_{n}$, or strategically with some a priori information on $\hat{P}_{n}$.
@@ V.1 Double-bracket iterations
$f\left(\hat{P}\right)\coloneq\lVert\hat{P}-\hat{D}\rVert_{2}^{2}\ .$ (S141)
The generator $\frac{d}{ds_{n}}\hat{P}_{n+1}=[[\hat{D},\hat{P}_{n}],\hat{P}_{n}]$ is shown to be $-\mathrm{grad}f(\hat{P}_{n})$ for some Riemannian metric defined on the manifold of matrices isospectral to $\hat{P}_{0}$ [47]. Hence, the iteration effectively applies the gradient flow of $f$ and we expect $f$ to reduce over the iterations towards the minimum. This is in fact true for suitable choices of $s_{n}$ and $\hat{D}$.
@@
When $\hat{D}$ is a non-degenerate diagonal matrix such that $\hat{D}=\sum_{i}\mu_{i}|{i}\rangle\!\langle{i}|$ with $\mu_{i}>\mu_{j}$ for all $i>j$ and $s_{n}=\frac{1}{4\lVert\hat{P}_{0}\rVert_{2}\lVert\hat{D}\rVert_{2}}$, the recursion Eq. (S140) has a unique stable fixed-point $\hat{P}_{\infty}=\sum_{i}\lambda_{i}|{i}\rangle\!\langle{i}|$, where $\lambda_{i}$ are eigenvalues of $\hat{P}_{0}$ arranged in a non-increasing order. Furthermore, this fixed-point is locally exponentially stable.
In particular, the off-diagonal element of $\hat{P}_{k}$ is suppressed after each step
$(\hat{P}_{k+1})_{ij}=\left[1-\frac{(\lambda_{i}-\lambda_{j})(\mu_{i}-\mu_{j})}{4\lVert\hat{P}_{0}\rVert_{2}\lVert\hat{D}\rVert_{2}}\right](\hat{P}_{k})_{ij}\ ,$ (S142)
up to the first order of $(\hat{P}_{k})_{ij}$, where the suppression factor inside the square bracket is in the range $(0,1)$. Hence, it is possible to establish the exponential suppression of the distance.
@@
Let $\hat{P}_{\infty}$ be a fixed-point of the double-bracket iterations starting from a matrix $\hat{P}_{0}$ and let $M(\hat{P}_{0})$ be a manifold of matrices isospectral to $\hat{P}_{0}$. Then there exists $r<1$ and a neighbourhood $N$ of $\hat{P}_{\infty}$ in the manifold $M(\hat{P}_{0})$, such that
$\lVert\hat{P}_{k+1}-\hat{P}_{\infty}\rVert_{2}<r\lVert\hat{P}_{k}-\hat{P}_{\infty}\rVert_{2}\ ,$ (S143)
for any $\hat{P}_{k}\in N$ and $\hat{P}_{k+1}$ obtained from an iteration from $\hat{P}_{k}$.
@@ V.2 Unfolding implementation for double-bracket iterations with black box queries
We now describe the unfolding implementation of double-bracket iterations, given access to the black box queries to the evolution $e^{\pm it\hat{P}_{0}}$ by the root operator $\hat{P}_{0}$, following Ref. [6]. To implement the memory-call $e^{i\mathcal{N}(\hat{P})}=e^{s_{n}[\hat{D}_{n},\hat{P}]}$ with queries $e^{\pm it\hat{P}_{0}}$, the group commutator approximation is used.
@@
$\left\lVert e^{-i\sqrt{s}\hat{B}}e^{-i\sqrt{s}\hat{A}}e^{i\sqrt{s}\hat{B}}e^{i\sqrt{s}\hat{A}}-e^{s[\hat{A},\hat{B}]}\right\rVert\leq s^{3/2}\left(\left\lVert[\hat{A},[\hat{A},\hat{B}]\right\rVert+\left\lVert[\hat{B},[\hat{B},\hat{A}]\right\rVert\right)\ .$ (S144)
Setting $\hat{A}=\hat{P}$ and $\hat{B}=-\hat{D}$ we obtain an locally accurate approximation to the recursion step
$\hat{W}_{n}=\left(e^{i\sqrt{\frac{s_{n}}{M}}\hat{D}_{n}}e^{-i\sqrt{\frac{s_{n}}{M}}\hat{P}_{n}}e^{-i\sqrt{\frac{s_{n}}{M}}\hat{D}_{n}}e^{i\sqrt{\frac{s_{n}}{M}}\hat{P}_{n}}\right)^{M}$ (S145)
that approximates $e^{s_{n}[\hat{D},\hat{P}_{n}]}$ with error $\mathcal{O}(s_{n}^{3/2}M^{-1/2})$ by making $L=2M$ calls to $P_{n}$. Instead of Eq. (S145), one can employ higher order product formulae [60] that can be more complicated to execute but have higher efficiency. Typically, we apply this recursion to a state
$\rho_{n}=\frac{\hat{P}_{n}+\lambda\mathmybb{1}}{\operatorname{Tr}\left[\hat{P}_{n}+\lambda\mathmybb{1}\right]}\ ,$ (S146)
encoding the matrix $\hat{P}_{n}$. The convergence of the recursion $\rho_{k+1}=\hat{W}_{k}\rho_{k}\hat{W}_{k}^{\dagger}$ to the diagonal state $\rho_{\infty}$ has been established, despite the error arising from the group commutator approximation [6].
The memory-call to $\hat{P}_{n}$ needed for Eq. (S145) is not directly accessible to us. However, it is covariant, i.e. $e^{-it\hat{P}_{n}}=\hat{W}_{n-1}e^{-it\hat{P}_{n-1}}\hat{W}_{n-1}^{\dagger}$. Unfolding then follows from replacing all memory-calls in Eq. (S145) by the call to the root state $e^{\pm it\hat{P}_{0}}$, dressed with previous recursion steps. The circuit depth scaling can also be calculated for the unfolding implementation. Since $L=2M$, the depth required for performing $N$ steps of the iteration would scale as $\mathcal{O}((4M+1)^{N})$.